Hi I'm back. Thank you so much for your help. However I'm having problems plugging in the numbers into the COS formula.
Find the area of a parallelogram whose sides are 6 inches and 8 inches and whose longer diagonal is 12 inches.
Could you show me how to plug in the numbers and find the area. Thank you so much.

Answer by venugopalramana(3286)

Find the area of a parallelogram whose sides are 6 inches and 8 inches and whose longer diagonal is 12 inches.
IF ABCD IS PARALLELOGRAM ,IT IS SPLIT INTO 2 EQUAL SIZED TRIANGLES BY DIAGONL SAY AC...SO AREA OF PARALLELOGRAM IS SUM OF AREAS OF 2 TRIANGLES WHICH ARE IDENTICAL IN AREA.SO WE NEED TO FIND AREA OF TRIANGLE ABC AND MULTIPLY IT WITH 2
NOW AREA OF ABC GIVEN ALL 3 SIDES
a=8...b=6...c=12..IS GIVEN BY
A=SQUARE ROOT OF {s(s-a)(s-b)(S-c)} WHERE S=SEMIPERIMETER=(a+b+c)/2=(8+6+12)/2=13
I THINK NOW YOU CAN FIND THE ANSWER AS YOU WANTED ONLY THE FORMULA